the measure of angle B is 20 degrees more than its compliment what is the measure of angle B the equation is x+20+x=90 which becomes 2x+20=90 which becomes 2/70=35 i just don't know what to do with angle B=x+20___degrees
Answer by Boreal(15235)

Complement of B is 90-B
B-20=90-B
add B to both sides
2B-20=90; add 20 to both sides
2B=110; divide both sides by 2
B=55 degrees
Complement is 35